# https://kartik-3004.github.io/facexformer_web/

### v1

@@ Abstract
Figure 1: FaceXformer an end-to-end unified transformer model for 8 different facial analysis tasks such as face parsing, landmark detection, head pose estimation, attributes recognition, and estimation of age, gender, race, and landmarks visibility.
In this work, we introduce FaceXformer, an end-to-end unified transformer model for a comprehensive range of facial analysis tasks such as face parsing, landmark detection, head pose estimation, attributes recognition, and estimation of age, gender, race, and landmarks visibility. Conventional methods in face analysis have often relied on task-specific designs and preprocessing techniques, which limit their approach to a unified architecture. Unlike these conventional methods, our FaceXformer leverages a transformer-based encoder-decoder architecture where each task is treated as a learnable token, enabling the integration of multiple tasks within a single framework. Moreover, we propose a parameter-efficient decoder, FaceX, which jointly processes face and task tokens, thereby learning generalized and robust face representations across different tasks. To the best of our knowledge, this is the first work to propose a single model capable of handling all these facial analysis tasks using transformers. We conducted a comprehensive analysis of effective backbones for unified face task processing and evaluated different task queries and the synergy between them. We conduct experiments against state-of-the-art specialized models and previous multi-task models in both intra-dataset and cross-dataset evaluations across multiple benchmarks. Additionally, our model effectively handles images "in-the-wild," demonstrating its robustness and generalizability across eight different tasks, all while maintaining the real-time performance of 37 FPS.
@@ 1 Introduction
Face analysis is a crucial problem as it has broad range of application such as ace verification and identification [61, 62], surveillance [19], face swapping [11], face editing [92], de-occlusion [79], 3D face reconstruction [75], retail [1], image generation [78] and face retrieval [81]. Facial analysis tasks (Figure 1 involve face parsing [25, 73], landmarks detection [39, 90], head pose estimation [89, 10], facial attributes recognition [48, 46], age/gender/race estimation [6, 37] and landmarks visibility prediction [43, 31]. Therefore, developing a generalized and robust face model for all tasks is a crucial and longstanding problem in the face community.
In recent years, significant advancements have been made in facial analysis, developing state-of-the-art methods for various tasks  [89, 90, 37, 10, 79, 11]. Despite these methods achieving promising performance, they cannot be integrated into a single pipeline due to their specialized model designs and task-specific pre-processing techniques. This limitation hinders the expansion toward developing a single unified model capable of solving all facial analysis tasks (see Table.1). A single model capable of addressing multiple facial tasks is desirable because it: (1) learns a robust and generalized face representation capable of handling in-the-wild images; (2) can model the intra-task relationship (e.g., the relationship between face parsing and landmark detection); and (3) improves task synergy leading to overall performance enhancement. In this work, we introduce FaceXformer, an end-to-end unified model designed for a comprehensive range of facial analysis tasks, such as face parsing, landmark detection, head pose estimation, attributes recognition, age/gender/race estimation and landmarks visibility prediction.
Methods FP LD HPE Attr Age Gen Race Vis MD Task Queries Shared Model
@@
Table 1: Comparison with representative methods under different task settings. Our proposed FacexFormer can perform various facial analysis tasks in single model. FP - Face Parsing, LD - Landmarks Detection, HPE - Head Pose Estimation, Attr - Attributes Recognition, Age - Age Estimatin, Gen - Gender Estimation, Race - Race Estimation, Vis - Landmarks Visibility Prediction, MD - Multi-dataset Training
FaceXformer enables task unification by leveraging the transformer as its core component. Specifically, we introduce a transformer-based encoder-decoder structure, treating each facial analysis task as a unique, learnable token within the framework. Treating each task as a token allows for the simultaneous processing of multiple facial analysis tasks, overcoming the challenges present in conventional methods that depend on separate, task-specific models and preprocessing routines. Furthermore, we introduce a parameter-efficient decoder, FaceX, which processes both face and task tokens together, enabling the model to learn robust face representations that generalize across various tasks. After modeling the intra-task and face-token relationships in the FaceX decoder, the task tokens are fed into a unified head, which essentially converts these task tokens into corresponding task predictions. We conduct experiments against state-of-the-art specialized models, and previous multi-task models, including both intra-dataset and cross-dataset evaluations across multiple benchmarks. Additionally, we performed extensive experiments to evaluate different task queries and the synergy between them and further experimented with the backbone to assess its effectiveness towards a unified model. Moreover, we show that our model effectively handles images ’in the wild,’ demonstrating its robustness and generalizability across eight different tasks. To the best of our knowledge, this is the first work to propose a single model capable of handling all these facial analysis tasks using transformers, all while maintaining real-time performance at 37 FPS. In summary, our paper’s contributions are as follows:
@@
    We introduce FaceXformer, a unified transformer-based framework for the simultaneous processing of multiple facial analysis tasks, eliminating the need for separate, task-specific models.
@@
    We propose the FaceX decoder, a parameter-efficient decoder that treats each facial analysis task as a token and processes them together, leading to improved generalization and task synergy across various tasks.
@@
    We conducted extensive experiments and analysis, including both intra-dataset and cross-dataset evaluations, comparing our approach to existing state-of-the-art specialized models and multi-task models across all eight tasks.
@@ 2 Related Work
Facial analysis tasks: Facial analysis tasks involve face parsing [25, 73, 9, 65, 87], landmarks detection [39, 33, 36, 90, 44], head pose estimation [89, 67, 82, 10], facial attributes recognition [48, 46, 58, 88], age/gender/race estimation [6, 32, 35, 37] and landmarks visibility prediction [43, 31]. These tasks hold significance in various applications such as face swapping [11], face editing [92], de-occlusion [79], 3D face reconstruction [75], driver assistance [47], human-robot interaction [59], retail [1], face verification and identification [61, 62],image generation [78], image retrieval [81] and surveillance [19]. Specialized models excel in their respective tasks but cannot be easily integrated with other tasks due to the need for extensive task-specific pre-processing [40, 90]. Generally, these models under-perform when applied to tasks beyond their specialization as their design is specific to their designated tasks. Some works [84, 45, 86, 24] perform multiple tasks simultaneously but utilize the additional tasks for guidance or auxiliary loss calculation to enhance the performance of the primary task. HyperFace [51] and AllinOne [52] are the only studies that explore multi-task learning with the aim of performing multiple tasks, but tend to limit their scope to tasks that inherently complement each other. For instance, landmark detection and head pose estimation are both regression-based tasks that effectively support each other. HyperFace utilizes multi-scale features from different layers of CNNs, and is capable of landmarks detection, head pose estimation and gender estimation, while AllinOne additionally performs face recognition and age estimation. Both these works, misleadingly suggest that they perform face detection, however, they use the selective search [55] algorithm in R-CNN [20] for region proposal, and merely predict whether the selected region is a face.
@@
Unified transformer models: In recent years, the rise of transformers [68, 15] have paved the way for the unification of multiple tasks within a single architecture. Unified transformer architectures are being explored across various computer vision problems, including segmentation [38, 98], visual question answering (VQA) [70, 80], tracking [71, 91], detection [72]. While these models may not achieve state-of-the-art (SOTA) performance and may under-perform compared to specialized models, they demonstrate satisfactory performance across a variety of tasks. Such unification efforts have led to the development of foundational models like SAM [30], CLIP [49], LLaMA [66], GPT-3 [4], DALL-E [50], etc. However, these models are computationally intensive and not suitable for facial analysis applications that require real-time performance. Motivated by this challenge, we propose FaceXFormer: the first lightweight, transformer-based model capable of performing multiple facial analysis tasks. FaceXFormer delivers real-time performance at $37$ FPS and can be integrated with existing face detectors to provide additional insights about the person of interest.
3 Proposed Work
@@
3.1 FaceXFormer
In our framework, we follow a standard encoder-decoder structure as illustrated in Fig. 2. For an input face image $\mathbf{I}\in\mathbb{R}^{H\times W\times 3}$, we extract coarse to fine-grained multi-scale features $\mathbf{S}_{i}$, where $i$ belongs to the $i$-th encoder output. To learn a unified face representation $\mathbf{F}$, these multi-scale features are then fused using a lightweight MLP-Fusion $\mathbf{M}$ module. Following fusion, we initialize a series of task-specific tokens $\mathbf{T}=\langle T_{1},\dots,T_{n}\rangle$, with each $t_{i}$ representing a face task. Afterward, we initialize task tokens $\mathbf{T}=\langle T_{1},\dots,T_{n}\rangle$, where $T_{i}$ denotes each task. Face tokens $\mathbf{F}$ and task tokens $\mathbf{T}$ are then processed by a Parameter-efficient Decoder $\mathbf{FXDec}$ where the task tokens are attended with face tokens to learn relevant task representation.
$\langle\mathbf{\hat{T}}\rangle=\mathbf{FXDec}\left(\langle\mathbf{F},\mathbf{T}\rangle;\mathbf{S}_{i}\right)$ (1)
where, $\mathbf{\hat{T}}$ is the output task tokens. Later these output task tokens are fed into unified heads, where each task token is refined and passed to the task head for corresponding task prediction.
@@ 3.2 Multi-scale Encoder
In the encoder, we employ a multi-scale encoding strategy to address the varying feature requirements intrinsic to each face analysis task. For instance, age estimation requires a global representation, while face parsing necessitates a fine-grained representation. Given an input image $\mathbf{I}$, it is processed through a set of encoder layers. For each encoder layer, the output captures information at varying levels of abstraction and detail, generating multi-scale features $\{\mathbf{S}_{i}\}_{i=1}^{n}$, where $i$ ranges from 1 to 4. This results in a hierarchical structure of features, wherein each feature map $\mathbf{S}_{i}$ transitions from a coarse to a fine-grained representation suitable for diverse facial tasks.
@@
where $D_{i}$ and $D_{t}$ are the multi-scale feature channel dimensions of $\mathbf{S}_{i}$ and the target channel dimension, respectively. The Lightweight MLP-Fusion design ensures minimal computational overhead while maintaining the ability to perform efficient feature fusion, crucial for subsequent face analysis tasks.
@@ 3.3 FaceX Decoder
$\mathbf{T}^{\prime}_{i}=\text{SelfAttn}(\mathbf{Q}=T^{\prime}_{i},\mathbf{K}=\mathbf{T},\mathbf{V}=\mathbf{T}),$ (2)
where Attention denotes the multi-headed self-attention mechanism, and $\mathbf{Q}$, $\mathbf{K}$, and $\mathbf{V}$ represent the queries, keys, and values, respectively. Therefore, TSA essentially helps to model the inter-task relationship and improve overall task synergy.
@@
$\hat{T}_{i}=\text{CrossAttn}(\mathbf{Q}={T}^{\prime}_{i},\mathbf{K}=\mathbf{F},\mathbf{V}=\mathbf{F}),$ (3)
@@
$\mathbf{\hat{F}}=\text{CrossAttn}(\mathbf{Q}=\mathbf{F},\mathbf{K}=\mathbf{T}^{\prime},\mathbf{V}=\mathbf{T}^{\prime}).$ (4)
Through this inverse attention mechanism, the face representation is augmented with critical task-specific details, enabling a more robust approach towards facial task unification.
@@ 3.4 Unified-Head
In Unified-Head, the task tokens are processed to obtain corresponding task predictions. As shown in Fig. 2, the output face tokens $\mathbf{\hat{F}}$ and task tokens $\mathbf{\hat{T}}$ are processed through a Task-to-Face Cross-Attention mechanism to obtain final refined features. Then, the output tokens are fed into their corresponding task heads. The task head for landmark detection and head pose estimation is a regression MLP, while the tasks of estimating age, gender, race, visibility, and attributes recognition utilize classification MLPs. For face parsing, we leverage the output $\mathbf{\hat{F}}$ and process it through an upsampling layer, then perform a cross-product with the face parsing token to obtain a segmentation map. Note, the number of tokens for segmentation corresponds to the total number of classes, and for regression and classification tasks, one token is used per task.
3.5 Training and Inference
@@
$L=\lambda_{seg}L_{seg}$ $+\lambda_{lnd}L_{lnd}$
$+\lambda_{hpe}L_{hpe}+\lambda_{attr}L_{attr}+\lambda_{a/g/r}L_{a/g/r}+\lambda_{vis}L_{vis},$ (5)
where $L_{seg}$ is the mean of dice loss [60] and Cross-Entropy (CE) loss for face parsing, $L_{lnd}$ is wing loss [18] for landmarks estimation, $L_{hpe}$ is geodesic loss [54] for head pose estimation, $L_{a/g/r}$ is CE loss for age/gender/race estimation, and $L_{attr}$ and $L_{vis}$ are Binary Cross-Entropy with logits loss for attributes recognition and landmarks visibility prediction respectively. During inference, FaceXFormer can be prompted for multiple outputs with task-specfic tokens. For instance, when prompting for parsing, head pose and age/gender/race, it generates outputs exclusively for those tasks, as illustrated in Figure 3.
Figure 3: Task-specific inference based on task-tokens on in-the-wild images. In the second column, we query the face parsing token. In the third column, we query the head pose token. In the fourth column, we query the infer age, gender, and race task tokens.

### v2

# FaceXFormer: A Unified Transformer for Facial Analysis

@@ Abstract
In this work, we introduce FaceXFormer, an end-to-end unified transformer model capable of performing nine facial analysis tasks including face parsing, landmark detection, head pose estimation, attribute prediction, and estimation of age, gender, race, expression, and face visibility within a single framework. Conventional methods in face analysis have often relied on task-specific designs and pre-processing techniques, which limit their scalability and integration into a unified architecture. Unlike these conventional methods, FaceXFormer leverages a transformer-based encoder-decoder architecture where each task is treated as a learnable token, enabling the seamless integration and simultaneous processing of multiple tasks within a single framework. Moreover, we propose a novel parameter-efficient decoder, FaceX, which jointly processes face and task tokens, thereby learning generalized and robust face representations across different tasks. We jointly trained FaceXFormer on nine face perception datasets and conducted experiments against specialized and multi-task models in both intra-dataset and cross-dataset evaluations across multiple benchmarks, showcasing state-of-the-art or competitive performance. Further, we performed a comprehensive analysis of different backbones for unified face task processing and evaluated our model “in-the-wild”, demonstrating its robustness and generalizability. To the best of our knowledge, this is the first work to propose a single model capable of handling nine facial analysis tasks while maintaining real-time performance at $33.21$ FPS.
Figure 1: FaceXFormer an end-to-end unified transformer model for 9 different facial analysis tasks such as face parsing, landmark detection, head pose estimation, attributes recognition, and estimation of age, gender, race, expression and face visibility.
@@ 1 Introduction
Face analysis is a crucial problem as it has broad range of application such as face verification and identification [84, 85], surveillance [22], face swapping [13], face editing [127], de-occlusion [111], 3D face reconstruction [102], retail [1], image generation [109] and face retrieval [113]. Facial analysis tasks (Figure 1 involve face parsing [28, 98], landmarks detection [46, 125], head pose estimation [124, 12], facial attributes recognition [64, 57], age/gender/race/expression estimation [8, 43] and landmarks visibility prediction [54, 35]. Therefore, developing a generalized and robust face model for all tasks is a crucial and longstanding problem in the face community.
In recent years, significant advancements have been made in facial analysis, developing state-of-the-art methods and face libraries for various tasks  [124, 125, 43, 12, 111, 13]. Despite these methods achieving promising performance, they cannot be integrated into a single pipeline due to their specialized model designs and task-specific pre-processing techniques. Furthermore, deploying multiple specialized models simultaneously is computationally intensive and impractical for real-time applications, leading to increased system complexity and resource consumption. These challenges emphasis the need for a unified model that can concurrently handle multiple facial analysis tasks efficiently (see Table 1). A single model capable of addressing multiple facial tasks is desirable because it: (1) learns a robust and generalized face representation capable of handling in-the-wild images; (2) intra-task modeling helps the models to learn task-invariant representation; and (3) simplifies deployment pipelines by reducing computational overhead and achieving real-time performance.
Methods FP LD HPE Attr Age Gen Race Vis Exp
Single-Task Models
@@
SSPL [80]
VOLO-D1 [36]
DLDL-v2 [21]
3DDE [89]
MNN [90]
KTN [40]
DMUE [77]
Multi-Task Models
@@
Table 1: Comparison with representative methods under different task settings. Our proposed FaceXFormer can perform various facial analysis tasks in a single model. FP - Face Parsing, LD - Landmarks Detection, HPE - Head Pose Estimation, Attr - Attributes Recognition, Age - Age, Gen - Gender, Race - Race Estimation, Exp - Facial Expression Recognition, and Vis - Face Visibility
To this end, we introduce FaceXFormer, an end-to-end unified model designed for nine different facial analysis tasks, such as face parsing, landmark detection, head pose estimation, attribute recognition, age/gender/race/expression estimation and face visibility prediction. FaceXFormer enables task unification by leveraging the transformers and learnable tokens as its core components. Specifically, we introduce a transformer-based encoder-decoder structure, treating each facial analysis task as a unique, learnable token within the framework. Treating each task as a token allows for the simultaneous processing of multiple facial analysis tasks, overcoming the challenges present in conventional methods that depend on separate, task-specific models and pre-processing routines. Furthermore, we introduce a parameter-efficient decoder, FaceX, which processes both face and task tokens together, enabling the model to learn robust face representations that generalize across various tasks. This parameter-efficient design reduces computational load and allows our model to perform in real-time. After modeling the intra-task and face-token relationships in the FaceX decoder, the task tokens are fed into a unified head, which essentially converts these task tokens into corresponding task predictions.
Our extensive experiments demonstrate that FaceXFormer achieves state-of-the-art or competitive performance compared to specialized models and existing multi-task frameworks, including both intra-dataset and cross-dataset evaluations across multiple benchmarks. Specifically, recent multi-task frameworks such as Faceptor [67], QFace [83], and Swinface [66] address fewer tasks than FaceXFormer, yet our model outperforms them on existing benchmarks. Moreover, we show that our model effectively handles images ’in the wild,’ demonstrating its robustness and generalizability across nine different tasks. This robustness is critical for real-world applications where uncontrolled conditions and diverse inputs are common. Achieving real-time performance at $33.21$ FPS marks a significant advancement, making FaceXFormer highly suitable for time-sensitive applications. To the best of our knowledge, this is the first work to propose a single model capable of handling nine different facial analysis tasks using transformers, all while maintaining real-time performance.
In summary, our paper’s contributions are as follows:
@@
    We introduce FaceXFormer, a unified transformer-based framework capable of simultaneously processing nine different facial analysis tasks, achieving real-time performance of $33.21$ FPS.
@@
    We propose FaceX, a parameter-efficient decoder that represents each facial analysis task as a token, enabling joint processing of face and task tokens.
@@
    We conduct extensive experiments and analyses, including both intra-dataset and cross-dataset evaluations, demonstrating that our approach achieves state-of-the-art performance when compared to existing specialized and multi-task models across multiple tasks.
@@ 2 Related Work
Facial analysis tasks: Facial analysis tasks involve face parsing [28, 98, 11, 86, 122, 62], landmarks detection [46, 37, 41, 125, 55], head pose estimation [124, 90, 115, 12], facial attributes recognition [64, 57, 80, 123], age/gender/race estimation [8, 36, 39, 43] and landmarks visibility prediction [54, 35]. These tasks hold significance in various applications such as face swapping [13, 60], face editing [127], de-occlusion [111], 3D face reconstruction [102], driver assistance [59], human-robot interaction [81], retail [1], face verification and identification [84, 85],image generation [109], image retrieval [113] and surveillance [22, 61]. Specialized models excel in their respective tasks but cannot be easily integrated with other tasks due to the need for extensive task-specific pre-processing [47, 125]. Generally, these models under-perform when applied to tasks beyond their specialization as their design is specific to their designated tasks. Some works [119, 56, 121, 26] perform multiple tasks simultaneously but utilize the additional tasks for guidance or auxiliary loss calculation to enhance the performance of the primary task. HyperFace [70] and AllinOne [71] are early models that explore multi-task learning with the aim of performing multiple tasks. HyperFace utilizes multi-scale features from different layers of CNNs, and is capable of landmarks detection, head pose estimation and gender estimation, while AllinOne additionally performs face recognition and age estimation. Recent models, such as SwinFace [66], QFace [83], and Faceptor [67], leverage transformer models with learnable tokens to perform multi-task learning. However, they perform fewer tasks, neglecting more complex ones such as head pose estimation, landmark estimation, and face parsing. The proposed FaceXFormer addresses all these tasks, alongside others, and achieves state-of-the-art performance in majority of them.
@@
Unified transformer models: In recent years, the rise of transformers [91, 17] have paved the way for the unification of multiple tasks within a single architecture. Unified transformer architectures are being explored across various computer vision problems, including segmentation [44, 132], visual question answering (VQA) [93, 112], tracking [96, 126], detection [97]. While these models may not achieve state-of-the-art (SOTA) performance and may under-perform compared to specialized models on some tasks, they demonstrate competitive performance across a variety of tasks. Such unification efforts have led to the development of foundational models like SAM [34], CLIP [68], LLaMA [88], GPT-3 [6], DALL-E [69], etc. However, these models are computationally intensive and not suitable for facial analysis applications that require real-time performance. Motivated by this challenge, we propose FaceXFormer: the first lightweight, transformer-based model capable of performing multiple facial analysis tasks. It delivers real-time performance at $33.21$ FPS and can be seamlessly integrated into existing systems providing additional annotations for the person of interest.
3 FaceXFormer
@@
In our framework, we follow a standard encoder-decoder structure as illustrated in Fig. 2. For an input face image $\mathbf{I}\in\mathbb{R}^{H\times W\times 3}$, we extract coarse to fine-grained multi-scale features $\mathbf{S}_{i}$, where $i$ belongs to the $i$-th encoder output. To learn a unified face representation $\mathbf{F}$, these multi-scale features are then fused using a lightweight MLP-Fusion $\mathbf{M}$ module. Following fusion, we initialize a series of task-specific tokens $\mathbf{T}=\langle T_{1},\dots,T_{n}\rangle$, with each $t_{i}$ representing a face task. Afterward, we initialize task tokens $\mathbf{T}=\langle T_{1},\dots,T_{n}\rangle$, where $T_{i}$ denotes each task. Face tokens $\mathbf{F}$ and task tokens $\mathbf{T}$ are then processed by a Parameter-efficient Decoder $\mathbf{FXDec}$ where task tokens are attended with face tokens to learn relevant task representation.
$\langle\mathbf{\hat{T}}\rangle=\mathbf{FXDec}\left(\langle\mathbf{F},\mathbf{T}\rangle;\mathbf{S}_{i}\right)$
Here, $\mathbf{\hat{T}}$ represents the output task tokens. These tokens are then fed into unified heads, where each task token is refined and passed to its respective task head for prediction.
@@ 3.1 Multi-scale Encoder
In the encoder, we employ a multi-scale encoding strategy to address the varying feature requirements intrinsic to each face analysis task. For instance, age estimation requires a global representation, while face parsing necessitates a fine-grained representation. Given an input image $\mathbf{I}$, it is processed through a set of encoder layers. For each encoder layer, the output captures information at varying levels of abstraction and detail, generating multi-scale features $\{\mathbf{S}_{i}\}_{i=1}^{n}$, where $i$ ranges from 1 to 4. This results in a hierarchical structure of features, wherein each feature map $\mathbf{S}_{i}$ transitions from a coarse to a fine-grained representation suitable for diverse facial analysis tasks.
@@
where $D_{i}$ and $D_{t}$ are the multi-scale feature channel dimensions of $\mathbf{S}_{i}$ and the target channel dimension, respectively. The lightweight MLP-fusion design ensures minimal computational overhead while maintaining the ability to perform efficient feature fusion, which is crucial for real-time application based face analysis tasks.
@@ 3.2 FaceX Decoder
$\mathbf{T}^{\prime}_{i}=\text{SelfAttn}(\mathbf{Q}=T^{\prime}_{i},\mathbf{K}=\mathbf{T},\mathbf{V}=\mathbf{T}),$
where Attention denotes the multi-headed self-attention mechanism, and $\mathbf{Q}$, $\mathbf{K}$, and $\mathbf{V}$ represent the queries, keys, and values, respectively. Therefore, TSA essentially helps the model to learn task-invariant representation.
@@
$\hat{T}_{i}=\text{CrossAttn}(\mathbf{Q}={T}^{\prime}_{i},\mathbf{K}=\mathbf{F},\mathbf{V}=\mathbf{F}),$
@@
$\mathbf{\hat{F}}=\text{CrossAttn}(\mathbf{Q}=\mathbf{F},\mathbf{K}=\mathbf{T}^{\prime},\mathbf{V}=\mathbf{T}^{\prime}).$
Through this inverse attention mechanism, the face representation is augmented with critical task-specific details, enabling a robust approach towards facial task unification.
@@ 3.3 Unified-Head
In Unified-Head, the task tokens are processed to obtain corresponding task predictions. As shown in Fig. 2, the output face tokens $\mathbf{\hat{F}}$ and task tokens $\mathbf{\hat{T}}$ are processed through a Task-to-Face Cross-Attention mechanism to obtain final refined features. Then, the output tokens are fed into their corresponding task heads. The task head for landmark detection is a hourglass network and for head pose estimation is a regression MLP, while the tasks of estimating age, gender, race, expression, visibility, and attributes prediction utilize classification MLPs. For face parsing, we leverage the output $\mathbf{\hat{F}}$ and process it through an upsampling layer, then perform a cross-product with the face parsing token to obtain a segmentation map. The number of tokens for segmentation corresponds to the total number of classes. For landmark prediction, it corresponds to the number of landmarks (i.e., $68$). For head pose estimation, the number of tokens is $9$, representing the $3\times 3$ rotation matrix. For other tasks, one token is used for each.
3.4 Multi-Task Training
@@
$L=\lambda_{seg}L_{seg}$ $+\lambda_{lnd}L_{lnd}+\lambda_{hpe}L_{hpe}+\lambda_{attr}L_{attr}$
$+\lambda_{a}L_{a}+\lambda_{g/r}L_{g/r}+\lambda_{exp}L_{exp}+\lambda_{vis}L_{vis}$
where $L_{seg}$ is the mean of dice loss [82] and Cross-Entropy (CE) loss for face parsing, $L_{lnd}$ is STAR loss [125] for landmarks prediction, $L_{hpe}$ is geodesic loss [115] for head pose estimation, $L_{g/r}$ is CE loss for gender/race estimation, $L_{a}$ is mean of L1 loss and CE loss for age estimation, $L_{exp}$ is CE loss for facial expression recognition, and $L_{attr}$ and $L_{vis}$ are Binary Cross-Entropy with logits loss for attributes prediction and face visibility prediction respectively.
Method Input Res. Skin Hair Nose L-Eye R-Eye L-Brow R-Brow L-Lip I-Mouth U-Lip Mean F1
Wei et al. [99] 512 96.4 91.1 91.9 87.1 85.0 80.8 82.5 91.0 90.6 87.9 88.43
EHANet [51] 512 96.0 93.9 93.7 86.2 86.5 83.2 83.1 90.3 93.8 88.6 89.53
EAGRNet [86] 473 96.2 94.9 94.0 88.6 89.0 85.7 85.2 91.2 95.0 88.9 90.87
AGRNet [87] 473 96.5 87.6 93.9 88.7 89.1 85.5 85.6 91.1 92.0 89.1 89.91
[…]
DML-CSR [122] 473 95.7 94.5 93.9 89.4 89.6 85.5 85.7 91.0 91.8 89.1 90.62
FP-LIIF [76] 256 96.4 95.1 93.7 88.5 88.5 84.5 84.3 90.3 92.1 87.5 90.09
SwinFace [66] $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
QFace [83] $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
Faceptor [67] 512 96.6 96.2 93.9 89.4 89.1 86.2 86.3 90.6 91.6 89.0 90.89
FaceXFormer 224 96.4 95.7 93.8 90.1 90.3 86.0 85.9 90.6 92.1 89.2 92.01
Table 2: Performance comparison for face parsing on the CelebAMask-HQ dataset [38]. The symbol $\times$ indicates that the model does not perform the corresponding task. Red = First Best, Blue = Second Best.